From a cylinder with a pressure of 60 bar and a volume of 40 l, a quantity of gas is withdrawn which, at an air pressure of 1 bar, has a volume of 80 l. Calculate the pressure in the bottle after the gas has been removed, if the temperature during the process remains constant
1
Expert's answer
2019-04-22T11:08:00-0400

 If the temperature during the process remains constant:


"pV=p'V+p_1V_1"

The pressure in the bottle after the gas has been removed:


"p'=p-p_1\frac{V_1}{V}"

"p'=60-1\frac{80}{40}=58 bar"